# Layerwise Recurrent Router for Mixture-of-Experts

### v1

@@ A.2 Experiment Setup
A.3 Additional Results
@@
A.4 Additional Observations
A.4.1 Mutual Information
1import numpy as np
2from sklearn.metrics import mutual_info_score
3
4def discretize_prob_dist(prob_dist, bins=100):
5 """
6 Discretize the probability distribution into discrete bins.
7 """
8 discretized = np.digitize(prob_dist, bins=np.linspace(0, 1, bins))
9 return discretized
10
11def calculate_mutual_information(x1, x2, bins=100):
12 """
13 Calculate mutual information between each pair of distributions in x1 and x2.
14 x1, x2: numpy arrays of shape (N, 16)
15 bins: number of bins to use for discretization
16 Returns a numpy array of mutual information values.
17 """
18 mi_values = []
19 for i in range(x1.shape[0]):
20 x1_discretized = discretize_prob_dist(x1[i], bins)
21 x2_discretized = discretize_prob_dist(x2[i], bins)
22 mi = mutual_info_score(x1_discretized, x2_discretized)
23 mi_values.append(mi)
24 return np.array(mi_values)
Refer to caption Refer to caption
Figure 7: Mutual information of RMoE-NP-r0.5 and CosineMoE settings
Refer to caption Refer to caption
Refer to caption Refer to caption
A.4.2 More Router Entropy Distributions
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to captionRefer to caption
Refer to caption Refer to caption
A.4.3 Expert Similarities
1def get_similarities(htoh4_0, htoh4_1, h4toh):
2 avg_key_0 = htoh4_0.mean(dim=1) # (num_experts, 4h, h)
3 avg_key_1 = htoh4_1.mean(dim=1) # (num_experts, 4h, h)
4 avg_value = h4toh.mean(dim=2) # (num_experts, h, 4h)
5 normed_key_0 = nn.functional.normalize(avg_key_0, p=2, dim=1)
6 normed_key_1 = nn.functional.normalize(avg_key_1, p=2, dim=1)
7 normed_value = nn.functional.normalize(avg_value, p=2, dim=1)
8 normed_avg_expert = torch.cat([normed_key_0, normed_key_1, normed_value], dim=1)
9 # compute the average expert similarity
10 similarity = torch.mm(normed_avg_expert, normed_avg_expert.t())
11 avg_sim = normed_similarity.mean().item()
12 return avg_sim
Refer to caption Refer to caption
A.4.4 Router Weights Information
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
A.4.5 Expert Selection Frequency
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption

### v2

@@ A.2 Experiment Setup
A.3 Further Pretraining Validation
To further validate the scalability of RMoE, we conduct experiments with larger model sizes and increased pre-training corpus. Both MoE models followed the design principles of DeepSeekMoE (Dai et al., 2024), utilizing fine-grained experts and shared experts to maintain strong baselines. We evaluated the models on more challenging benchmarks, including Hellaswag, MMLU, GSM8K, and HumanEval, to assess their language capabilities, multi-domain knowledge, mathematical skills, and coding abilities. Additionally, we tested the models’ perplexity on multiple domain test datasets and reported the average results.
Tab. 9 and Tab. 10 present the performance of a 15-billion parameter model with 2.7 billion activated experts, trained on 120 billion and 400 billion tokens, respectively. The results show that RMoE consistently delivers improvements even with increased data volumes. The findings indicate that RMoE enhances performance in standard language modeling tasks, such as Hellaswag and PPL, and improves performance on more complex reasoning tasks.
Table 9: Performance comparison of SMoE, SMoE-MLP and RMoE at the model scale of 15B activation 2.7B parameters, training 120B tokens.
Hellaswag MMLU GSM8K Avg PPL
Pretrain 80B Tokens
SMoE 67.69 46.24 24.18 7.406
SMoE-MLP 67.98 46.47 23.58 7.437
RMoE 68.00 47.74 27.14 7.361
Pretrain 100B Tokens
SMoE 70.98 50.61 30.78 6.754
SMoE-MLP 70.8 50.6 30.17 6.786
RMoE 71.02 51.74 32.98 6.732
Pretrain 120B Tokens
SMoE 72.03 52.79 34.8 6.447
SMoE-MLP 72.19 52.81 34.57 6.479
RMoE 72.36 54.02 36.13 6.425
Table 10: Performance comparison of SMoE, SMoE-MLP and RMoE at the model scale of 15B activation 2.7B parameters, training 400B tokens.
Hellaswag MMLU GSM8K Avg PPL
Pretrain 200B Tokens
SMoE 69.48 49.96 33.21 7.718
SMoE-MLP 69.76 50.27 31.77 7.736
RMoE 70.00 52.21 32.98 7.608
Pretrain 280B Tokens
SMoE 72.40 54.66 42.61 6.477
SMoE-MLP 72.62 55.33 38.51 6.502
RMoE 73.18 56.06 44.35 6.400
Pretrain 400B Tokens
SMoE 76.39 59.54 52.16 5.685
SMoE-MLP 76.09 59.96 51.71 5.709
RMoE 76.72 60.60 52.99 5.620
A.4 Additional Observations
A.4.1 Router Gradient Norm and Drop Ratio
Table 11: Comparison of linear and RNN routers in terms of gradients and drop ratios at various training steps. We record the router gradient every 10k training steps (20B tokens). We compute the gradient with language modeling (LM) loss and load balance (LB) loss. Drop ratio is the ratio of dropped tokens and all tokens as we assign capacity factor 1.0 for each expert.
Training steps (k step) 0.1 10 20 30 40 50 60
Linear router
grad from the whole loss 1.058 0.194 0.1911 0.198 0.208 0.217 0.221
grad from LM loss 0.625 0.183 0.184 0.192 0.204 0.215 0.220
grad from LB loss 0.433 0.011 0.008 0.006 0.004 0.002 0.001
drop ratio 35.6 5.43 5.34 5.17 4.89 4.64 4.50
RNN router
grad from the whole loss 0.972 0.160 0.153 0.153 0.155 0.155 0.154
grad from LM loss 0.636 0.146 0.138 0.139 0.144 0.148 0.151
grad from LB loss 0.337 0.014 0.015 0.014 0.011 0.007 0.003
drop ratio 38.7 6.35 6.30 5.94 5.32 4.54 4.09
Based on the setting of training 15B models for 120B tokens, we investigate how the gradient norm of the router varies throughout the entire training process. When training an MoE-based model, the gradient of the router has two separate sources: due to (1) the language modeling (LM) loss, and (2) the load balancing (LB) loss that forces the router to assign tokens to different experts in a balanced manner. Therefore, for each router, we compare the gradient from the LM loss only and from the whole training loss. We calculate the average for 100 training steps to estimate the gradient norm.
Furthermore, to better investigate the relation between the router behavior and the router gradient, we calculate the drop ratio for the router. This is because during the large-scale MoE pre-training, to ensure the training efficiency, the expert is usually controlled by an hyper-parameter called capacity factor, which determines the total tokens that one expert can process. If the router assigns tokens to some expert that exceeds its capacity, the expert will drop tokens with the lowest scores. And we define the drop ratio as tokens dropped / total tokens. The LB loss mentioned before is critical to decreasing the drop ratio.
According to Tab. 11, we have the following observations: 1. The gradient norm of the RNN router is generally smaller than that of the linear router. And for both routers, the drop ratio decreases with the training. 2. According to the drop ratios, we observe the significant behavioral difference between the two routers: during the early training phase (10k steps -¿ 30k steps), the drop ratio of the linear router is noticeably lower than that of the RNN router; the drop ratio of the RNN router archives at the lower value in the end. 3. The trend observed in the drop ratio is consistent with the results of the gradient norm. The grad norm for LB loss is relatively higher in the RNN router until the final training stage (50k - 60k), whereas the gradient from LB loss in the linear router is high at the beginning and generally low during the later part of training (10k - 60k).
These phenomena indicate that the LB loss could dominate the training of the linear router: when the drop ratio is low and stays unchanged, the grad from LB loss will be low because the router is already well-optimized for LB loss. Such early convergence in the LB loss may reach a suboptimal solution in the trade-off between optimizing load balance and language modeling. On the contrary, the gradient of the RNN router from LB loss stabilizes in the early training steps (10k - 30k), and the gradient from the lm loss keeps decreasing, suggesting that the RNN router is more optimized towards the LM loss.
A.4.2 Mutual Information
import numpy as np
from sklearn.metrics import mutual_info_score
def discretize_prob_dist(prob_dist, bins=100):
Discretize the probability distribution into discrete bins.
discretized = np.digitize(prob_dist, bins=np.linspace(0, 1, bins))
return discretized
def calculate_mutual_information(x1, x2, bins=100):
Calculate mutual information between each pair of distributions in x1 and x2.
x1, x2: numpy arrays of shape (N, 16)
bins: number of bins to use for discretization
Returns a numpy array of mutual information values.
mi_values = []
for i in range(x1.shape[0]):
x1_discretized = discretize_prob_dist(x1[i], bins)
x2_discretized = discretize_prob_dist(x2[i], bins)
mi = mutual_info_score(x1_discretized, x2_discretized)
mi_values.append(mi)
return np.array(mi_values)
A.4.3 Expert Similarities
def get_similarities(htoh4_0, htoh4_1, h4toh):
avg_key_0 = htoh4_0.mean(dim=1) # (num_experts, 4h, h)
avg_key_1 = htoh4_1.mean(dim=1) # (num_experts, 4h, h)
avg_value = h4toh.mean(dim=2) # (num_experts, h, 4h)
normed_key_0 = nn.functional.normalize(avg_key_0, p=2, dim=1)
normed_key_1 = nn.functional.normalize(avg_key_1, p=2, dim=1)
normed_value = nn.functional.normalize(avg_value, p=2, dim=1)
normed_avg_expert = torch.cat([normed_key_0, normed_key_1, normed_value], dim=1)
# compute the average expert similarity
similarity = torch.mm(normed_avg_expert, normed_avg_expert.t())
avg_sim = normed_similarity.mean().item()
return avg_sim
Refer to caption Refer to caption
Figure 6: Mutual information of RMoE-NP-r0.5 and CosineMoE settings
Refer to caption Refer to caption
Refer to caption Refer to caption
A.4.4 More Router Entropy Distributions
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to captionRefer to caption
Refer to caption Refer to caption
A.4.5 Router Weights Information
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
A.4.6 Expert Selection Frequency
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
A.5 Additional Results